10-2: Basic Definitions WACC Klose Outfitters Inc. believes that its optimal capital structure consists of 65% common equity and 35% debt, and its tax rate is 40%. Klose must raise additional capital to fund its upcoming expansion. The firm will have $2 millior of retained earnings with a cost of rs 15%. New common stock in an amour to $9 million would have a cost of re-: 17%. Furthermore, Klose can raise up to $3 million of debt at an interest rate of rd 9%, and an additional $4 nillion of debt at rd--12%. The CFO estimates that a proposed expansion ould require an investment of $5.0 million. What is the WACC for the last ollar raised to complete the expansion? Round your answer to two decimal laces. up
Answer #1
1) Component cost:
Cost of retained earnings (given) 15.00%
Cost of common equity (given) 17.00%
After tax cost of debt (upto debt of $3 million) = 9%*(1-40%) = 5.40%
After tax cost of debt (for further debt of $4 million) = 12%*(1-40%) = 7.20%
2) Debt required for $5 million = 5*35% = $         1.75 million
Equity required for $5 million = 5*65% = $         3.25 million
Retained earnings available $         2.00 million
New equity required = $3.25 million-$2.00 million = $         1.25 million
3) First stage Capital structure:
Retained earnings $         2.00 million
Debt = 2*35%/65% = $         1.08 million
Total $         3.08 million
Second stage capital structure:
Balance to be raised = 5-3.08 = $         1.92 million
Debt 35% and new equity 65%
WACC of last dollar raised = 5.40%*35%+17%*65% = 12.94%
Note:
Cost of debt upto $2 million is 5.40%. Total debt
upto capital requirement of $5 million is 1.75% only.
Hence, cost of debt is in the first slab; is 5.40%.